OpenStudy (anonymous):

Find the area of a regular hexagon with a base of 10 and an apothem of 20.

OpenStudy (anonymous):

360

OpenStudy (anonymous):

did this on my calculator, not sure about it, but: 1039.23048

OpenStudy (anonymous):

are you sure its not one of the choices

OpenStudy (anonymous):

desinitly not that but its 360

OpenStudy (anonymous):

the choices are 200 1200 600 300

OpenStudy (anonymous):

yeah im just trolling, i was gonna say 360, but she did already so yeah xD

OpenStudy (ccswims):

A = (1/2)ap, A = (1/2)(12)(60) = 360 units

OpenStudy (anonymous):

see i told u its 360

OpenStudy (anonymous):

your choices are weird, why isnt there 360?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

maybe u have to round, is this for homework?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

round it i guess, unless if there is something else that you didn't mention...

OpenStudy (anonymous):

thats wierd maybe you should ask yu teacher beacayse its 360 for sure

OpenStudy (anonymous):

Find the area of a regular hexagon with a base of 10 and an apothem of 20. 200 1200 600 300

OpenStudy (anonymous):

maybe you should just round

OpenStudy (anonymous):

how woud i round it

OpenStudy (anonymous):

round 360 to the closest huindred so whats the closest number to 360

OpenStudy (anonymous):

im not sure please help me

OpenStudy (anonymous):

300

OpenStudy (anonymous):

600 is too big 200 is to small and 300 is right

myininaya (myininaya):